# Literature Review, Methodology, Software, Data curation, Implementation, Experimental analysis, Writing - Original draft preparation

### v1

@@ Abstract
Phishing email is a serious cyber threat that tries to deceive users by sending false emails with the intention of stealing confidential information or causing financial harm. Attackers, often posing as trustworthy entities, exploit technological advancements and sophistication to make detection and prevention of phishing more challenging. Despite extensive academic research, phishing detection remains an ongoing and formidable challenge in the cybersecurity landscape. Large Language Models (LLMs) and Masked Language Models (MLMs) possess immense potential to offer innovative solutions to address long-standing challenges. In this research paper, we present an optimized, fine-tuned transformer-based DistilBERT model designed for the detection of phishing emails. In the detection process, we work with a phishing email dataset and utilize the preprocessing techniques to clean and solve the imbalance class issues. Through our experiments, we found that our model effectively achieves high accuracy, demonstrating its capability to perform well. Finally, we demonstrate our fine-tuned model using Explainable-AI (XAI) techniques such as Local Interpretable Model-Agnostic Explanations (LIME) and Transformer Interpret to explain how our model makes predictions in the context of text classification for phishing emails.
@@ 1 Introduction
The rapid advancement of internet technology has significantly changed the way people communicate online while also raising more security issues (Salloum et al., 2021). One of the prevalent security concerns, the cyberattack termed “email phishing” poses grave risks such as identity or login credentials theft, financial repercussions, and harm to an organization’s reputation. In this current decade, email phishing has changed throughout time to become a more effective method for carrying out fraudulent actions. As phishers continuously innovate new techniques, detecting phishing has become increasingly challenging, posing a formidable task for law enforcement agencies, governments, and organizations combating this cyber threat. To evade falling victim to this threat, organizations need to adeptly detect and prevent these attacks at scale by adopting effective phishing detection techniques. Leveraging Artificial Intelligence (AI) techniques like Machine Learning, Deep Learning, and Hybrid Learning which are already utilized for phishing attack detection, can significantly bolster these efforts (Basit et al., 2021) (Sarker, 2023). Moreover, because of the rapid advancements in technology, new methods based on AI are being developed daily in the cybersecurity field.
@@
Transformer-based models become more popular in the context of text classification. The transformer is an innovative neural design that uses the attention mechanism to encode incoming data into powerful characteristics (Han et al., 2021). Transformer-based models have greatly improved the processing, comprehension, and interpretation of text data inputs, which has had a revolutionary effect on the development of phishing classification techniques (Jamal et al., 2023). Large Language Models like OpenAI’s ChatGPT (Generative Pre-trained Transformer) and Bard have transformed the landscape of natural language creation and comprehension, capturing considerable attention from the general public (Zhao et al., 2023; Yao et al., 2023). The increasing complexity of LLMs is radically changing artificial intelligence and has the potential to transform algorithm development and application by understanding human language through extensive text data patterns and correlations.
@@
Many transformer-based models used in natural language processing (NLP), including well-known ones like OpenAI’s GPT and Google’s BERT (Bidirectional Encoder Representations from Transformers), are categorized as Large Language Models because of their ability to comprehend and produce text that is similar to that of a human. BERT model, a type of transformer-based neural network architecture pre-trained on extensive text corpora to grasp contextual word embeddings (Koroteev, 2021), become increasingly efficient and valuable in text classification and generation, including tasks like spam and phishing detection, sentiment analysis (Singh et al., 2021), and more other fields. BERT is designed to take into account both left and right contexts at every level to pre-train complex bidirectional representations from unlabeled text (Devlin et al., 2018). With simply one extra output layer, this method allows the pre-trained BERT model to be refined, leading to the development of state-of-the-art models for a variety of applications. In order to model language, BERT uses a Masked Language Model, in which certain words in a sentence are randomly replaced with a unique token (usually [MASK]). Next, by examining the contextual clues offered by nearby words, the model is trained to infer the original words. BERT makes use of a Transformer architecture technique called “self-attention”, which allows it to recognize inter-word relationships in text sequences and assess the relative importance of individual words within a sentence in order to understand the context. Based on the transformer layers number and parameters BERT model has different variants such as: BERT-base (Khadhraoui et al., 2022), ALBERT (Lan et al., 2019), RoBERTa (Liu et al., 2019), DistilBERT (Sanh et al., 2019), etc.
@@
Despite BERT’s versatility across NLP tasks, comprehending its internal mechanisms often poses challenges, resembling a “black box” due to its complexity. Unraveling how the model operates and forms predictions demands considerable effort (Xu et al., 2019). XAI techniques emerge as a vital solution, offering significant perceptions of the model’s reasoning and illuminating the decision-making process (Hoffman et al., 2018; Anan et al., 2023). Recently, Sarker (Sarker, 2024) has also highlighted the significance of explainability analysis in various perspectives in the context of cybersecurity, in which we are interested to explore more with experiments. XAI techniques can illuminate the words or features deemed most influential by the BERT model in specific text classification tasks of any nature. In this research, our main goal is to show transparency in phishing email detection with explainable AI. The major contribution of the work is as follows:
@@
  • We present a transformer-based Phishing email detection model, a fine-tuned version of DistilBERT.
  • Two different XAI techniques LIME and Transformer Interpret are utilized to explore the interpretability of the proposed fine-tuned model and visualize the scoring percentage of those techniques on prediction.
  • We demonstrate our model performance on both imbalanced and balanced datasets.
@@ 2 Literature Review
The field of cybersecurity has grown increasingly important and complex in the modern world. There are several subfields within cybersecurity, including data security, application security, network security, etc. Cyber-attacks of all kinds are becoming more and more frequent, and they can destroy or steal imported data. In addition, in this cybersecurity field, there are an enormous amount of new and creative techniques being developed daily for identifying and preventing cyberattacks. For the improvement of phishing email detection and prevention systems, many researchers have developed a wide range of novel strategies to counteract phishing emails, utilizing state-of-the-art methods from deep learning, machine learning, and hybrid models. Phishing is still a serious cybersecurity concern and one of the main ways hackers breach networks.
@@
Detecting phishing attempts early on is critical for businesses nowadays. The security of an organization can be significantly increased by machine learning, which can identify potential assaults early on (Apruzzese et al., 2023). In ref. (Yasin and Abuhasan, 2016), a text processing, data mining, and knowledge discovery approach was combined to create an intelligent phishing email detection model. They utilized text stemming and WordNet ontology in pre-processing to enrich the model with synonyms and employed five machine learning classification algorithms (J48, Naïve Bayes, Support Vector Machine (SVM), Multi-Layer Perceptron (MLP), and Random Forest (RF)). They have notably achieved 99.1% accuracy with RF and 98.4% using J48 for an accredited dataset. The study (Harikrishnan et al., 2018) employed techniques such as SVD and NMF for feature extraction and dimensionality reduction from legitimate and phishing email text data. For the numeric representation of phishing emails, the author integrated TFIDF with the SVD and NMF, and Decision Trees, Logistic Regression, Random Forest, Naive Bayes, K-Nearest Neighbors (KNN), AdaBoost, and SVM have been utilized as classifiers. Hamid et al. (2013) presented a hybrid feature selection technique combining content-based and behaviour-based approaches to detect phishing emails. Identifying the behaviour-based features in phishing emails was the main objective of this research because an attacker cannot disguise these types of features. They examined the overall process with a phishing dataset that consists of 6923 emails where phishing emails were collected from the Nazario dataset and legitimate emails from the SpamAssassin dataset and achieved 94% accuracy with this hybrid feature selection technique. With an emphasis on the importance of top-ranking features, the study (Zamir et al., 2020) examines feature selection methods (GR, IG, Relief-F, RFE) on a dataset of 11,055 phishing sites with 32 attributes. To improve classifier accuracy, it suggests new attributes that are generated from the strongest and weakest features. A variety of supervised learning methods with differing feature sets are compared, such as RF, SVM, bagging, kNN, Neural Network (NN), and j48. The proposed method using RF showed 97.3% and using stacking achieved 97.4% accuracy.
@@
In cybersecurity, deep learning has great potential, particularly for detecting network intrusions, classifying malware, and recognizing phishing attempts.Alhogail and Alsabih (2021) proposed a deep learning-based phishing email detection model based on graph convolutional network (GCN) and natural language processing (NLP). A publicly available balanced phishing dataset has been utilized where the authors used the email body’s text to extract features that improved the proposed model’s accuracy. The model showed 98.2% accuracy with a low false-positive rate of 0.015%. For phishing detection, a new Intelligent Cuckoo Search (CS) Optimization Algorithm with a Deep Learning-based Phishing Email Detection and Classification (ICSOA-DLPEC) approach has been proposed in (Brindha et al., 2023). Data preprocessing (email cleaning, tokenization, and stop-word elimination) techniques were employed to derive the important features from the Enron email dataset. The Gated Recurrent Unit (GRU) model was used with a CS algorithm for the fine-tuning of the parameter of the proposed model and achieved a high accuracy of 99.72%.Dewis and Viana (2022) employed Long short-term memory (LSTM) model with NLP to detect both phishing and spam email detection. They developed a phish responder which is a Python-based command line solution and used deep learning and NLP for this classification task. The proposed LSTM model achieved 99% accuracy for the text-based dataset. Fang et al. (2019) introduced THEMIS, a phishing email detection model built on an improved recurrent convolutional neural networks (RCNN) model which was improved by using the Bidirectional Long Short-Term Memory (Bi-LSTM). 99.84% accuracy is demonstrated by the suggested THEMIS model on datasets that are not balanced where the positive rate (FPR) is 0.043%. Other deep learning algorithms, such the reinforcement learning, Recurrent Neural Network (RNN), and Deep Belief Network (DBN), are also in use (Zhang and Li, 2017; Bahnsen et al., 2017; Smadi et al., 2018).
@@
Moreover, transformer-based models are also making a good impact in the cybersecurity field.Thapa et al. (2023) concentrated on two specific models: RNN and BERT to test how well these models could identify phishing emails. They looked at how well these models performed when the email data was distributed unevenly across several organizations and when added more data. In this experiment, the BERT model’s test accuracy was 96.1%. In addition, another study (Atawneh and Aljehani, 2023) focused on different deep learning architectures (CNNs, LSTMs, RNNs, BERT) to identify the emails as legitimate or phishing by analyzing different features of emails. The proposed model using BERT and LSTM, reaching an impressive accuracy of 99.61% on publicly available 4 different datasets, outperforming other methods. Nowadays, other variants of the BERT model such as DistilBERT, RoBERTa, and TinyBER (Jiao et al., 2019) are performing well in the phishing and spam detection research sector (Jamal et al., 2023; Lee et al., 2020; Songailaitė et al., 2023). The authors (Jamal et al., 2023) presented an Improved Phishing and Spam Detection Model (IPSDM) which consists of two BERT variants such as the fine-tuned DistilBERT and RoBERTA model and showed good accuracy in this research scope. Lee et al. Lee et al. (2020) proposed Context-Aware Tiny Bert (CatBERT) which exhibits faster processing and robustness to adversarial attacks than other transformer models, especially those that involve intentional keyword substitutions with synonyms or typos. The detection rate of this proposed model was 87% which was far better than other compared models. Furthermore, researchers also used fine-tuned BERT-based models for phishing URL detection (Wang et al., 2023; Maneriker et al., 2021).
@@
Overall, this literature survey shows the existing works although showing good performance by using different datasets and models but most of the research didn’t work model’s interpretability or explainability. “How does a model work?” or “Which things or materials contribute to detecting phishing?” those types of questions can arise in their working procedure. In this work, to answer those types of questions and solve the existing drawbacks in the literature, we have worked with explainable AI using not only one technique but also used two different techniques with different methodologies. In the literature, although some researchers did full experiments for phishing email detection but they utilized spam datasets as phishing datasets. In cybersecurity, we know that two different types of emails exist: spam and phishing, and not all spam emails are considered to be phishing. For this reason, we only focused on the phishing email dataset.
@@ 3 Methodology
In this paper, we have tried to explore the Transformer-based self-attention mechanism BERT model with the explainable AI. Initially, we collected a dataset and prepared it with data cleaning and balancing techniques. After preparing the data we have focused on optimizing and fine-tuning the model. For this phishing email detection purpose, we have selected the DistilBERT model and pre-trained with the prepared data. The DistilBERT model demonstrates good performance in both balanced and imbalanced scenarios of datasets when identifying phishing emails from texts. At the end of the experiment, we employed two XAI techniques Explanations LIME and Transformer Interpret to explore our fine-tuned model and described how it is working and performing with good accuracy in this text classification. The overall methodology of the proposed method is depicted in Fig. 1.
Figure 1: The proposed overall methodology.
3.1 Data Collection and Processing
For this experiment, a publicly available dataset has been collected from kaggle. The dataset describes different kinds of email text bodies that can be used to identify phishing emails using in-depth text analysis. This dataset consists of “Safe Email” and “Phishing Email”. Fig. 2 shows the data overview. The dataset has 11,322 Safe Emails and 7,328 Phishing Emails which shows that it is an imbalanced dataset. Along with this, this dataset showed that null values exist in rows. As null values can’t be processed that’s why those rows are removed. By using the resampling technique minor class “Phishing Email” is oversampled. This resampling technique mainly generates copies of other instances and adds those copies to the dataset. It is essential to address uneven data in order to develop equitable and precise predictive models. Balancing data ensures algorithms work well across all categories by reducing biases toward the majority class. This prevents skewed outputs and makes it possible to create a more accurate and equitable model for underrepresented classes. After oversampling, the dataset has 11,322 safe emails and 11,322 phishing emails in equal proportion. This data balancing also helps to reduce the risk of oversampling. Fig. 3 illustrates the imbalance and balanced data before and after sampling.
Figure 2: The sample data overview

### v2

# Literature Review, Methodology, Software, Data curation, Implementation, Experimental analysis, Conceptualization, Writing - Original draft preparation

@@ Abstract
Phishing email is a serious cyber threat that tries to deceive users by sending false emails with the intention of stealing confidential information or causing financial harm. Attackers, often posing as trustworthy entities, exploit technological advancements and sophistication to make the detection and prevention of phishing more challenging. Despite extensive academic research, phishing detection remains an ongoing and formidable challenge in the cybersecurity landscape. Large Language Models (LLMs) and Masked Language Models (MLMs) possess immense potential to offer innovative solutions to address long-standing challenges. In this research paper, we present a fine-tuned transformer-based language model, RoBERTa (Robustly Optimized BERT Pretraining Approach), for detecting phishing emails. In the detection process, we employ a real-world phishing email dataset and apply the preprocessing techniques to clean and address the class imbalance issues, thereby enhancing model performance. Our experimental results demonstrate that the fine-tuned model achieves high accuracy, confirming its robustness and effectiveness in the phishing email classification task. To ensure model transparency and trust in the model’s decision-making, we propose a hybrid explanation approach, LITA (LIME-Transformer Attribution), which integrates the potential of Local Interpretable Model-Agnostic Explanations (LIME) and Transformers Interpret methods. This hybrid method provides more user-friendly insights and solves the disagreement issues between the two approaches. Moreover, we also present how the model generates its predictions by presenting positive and negative contribution scores using LIME, Transformers Interpret, and LITA.
@@ 1 Introduction
The rapid advancement of internet technology has significantly changed the way people communicate online while also raising more security issues (Salloum et al., 2021). One of the prevalent security concerns, the cyberattack termed “email phishing” poses grave risks such as identity or login credentials theft, financial repercussions, and harm to an organization’s reputation (Almomani et al., 2013). In this current decade, email phishing has changed throughout time to become a more effective method for carrying out fraudulent actions. As phishers continuously innovate new techniques, detecting phishing has become increasingly challenging, posing a formidable task for law enforcement agencies, governments, and organizations combating this cyber threat. To evade falling victim to this threat, organizations need to adeptly detect and prevent these attacks at scale by adopting effective phishing detection techniques. Leveraging Artificial Intelligence (AI) techniques like Machine Learning, Deep Learning, and Hybrid Learning, which are already utilized for phishing attack detection, can significantly bolster these efforts (Basit et al., 2021) (Sarker, 2023). Moreover, because of the rapid advancements in technology, new methods based on AI are being developed daily in the cybersecurity field.
@@
Transformer-based models have become more popular in the context of text classification. The transformer is an innovative neural design that uses the attention mechanism to encode incoming data into powerful characteristics (Han et al., 2021). Transformer-based models have greatly improved the processing, comprehension, and interpretation of text data inputs, which has had a revolutionary effect on the development of phishing classification techniques (Jamal et al., 2023). LLMs like OpenAI’s ChatGPT (Generative Pre-trained Transformer) and Google’s Bard have transformed the landscape of natural language creation and comprehension, capturing considerable attention from the general public (Zhao et al., 2023; Yao et al., 2023). The increasing complexity of LLMs is radically changing artificial intelligence and has the potential to transform algorithm development and application by understanding human language through extensive text data patterns and correlations.
@@
Many transformer-based models used in natural language processing (NLP), including well-known ones like OpenAI’s GPT and Google’s BERT (Bidirectional Encoder Representations from Transformers), are categorized as language models due to their large number of parameters and are designed to process and generate human language. The BERT model, a type of transformer-based neural network architecture pre-trained on extensive text corpora to capture contextual word embeddings (Koroteev, 2021), has become increasingly efficient and valuable in text classification and generation tasks, including spam and phishing detection, sentiment analysis (Singh et al., 2021), and various other fields. BERT utilizes a bidirectional representation approach that considers both left and right context simultaneously at all layers, allowing it to learn rich language representations from large amounts of unlabeled text (Devlin et al., 2018). With simply one extra output layer, this method allows the pre-trained BERT model to be refined, leading to the development of state-of-the-art models for a variety of applications. In order to model language, BERT uses a Masked Language Model, in which certain words in a sentence are randomly replaced with a unique token (usually [MASK]). Next, by examining the contextual clues offered by nearby words, the model is trained to infer the original words. BERT makes use of a Transformer architecture technique called “self-attention”, which allows it to recognize inter-word relationships in text sequences and assess the relative importance of individual words within a sentence in order to understand the context. Based on the transformer layers number and parameters BERT model has different variants, such as: BERT-base (Khadhraoui et al., 2022), ALBERT (Lan et al., 2019), RoBERTa (Liu et al., 2019), DistilBERT (Sanh et al., 2019), etc.
@@
Despite BERT’s versatility across NLP tasks, comprehending its internal mechanisms often poses challenges, resembling a “black box” due to its complexity. Unraveling how the model operates and forms predictions demands considerable effort (Xu et al., 2019). XAI techniques emerge as a vital solution, offering significant perceptions of the model’s reasoning and illuminating the decision-making process (Hoffman et al., 2018; Anan et al., 2023). Recently, Sarker (Sarker, 2024; Sarker et al., 2024) has also highlighted the significance of explainability analysis in various perspectives in the context of cybersecurity, in which we are interested in exploring more with experiments. Explainable methods can highlight the words, sub-words, or features that the language model considers most influential in various text classification tasks. In our previous work (Uddin et al., 2025), we implemented two explainable methods and provided explanations with positive and negative coefficients of words. In this research, our main goal is to focus transparency in phishing email detection with explainable AI and to propose a hybrid method that considers the limitations of existing approaches. The major contributions of the work are as follows:
@@
  • We build a robust phishing email detection model by fine-tuning a pre-trained RoBERTa transformer-based language model.
  • We present a hybrid explanation approach, LITA, to provide a more user-friendly explanation for transformer-based phishing email detection. In addition, to produce alternate explanations, we also use LIME and Transformers Interpret, which enables us to compare results and draw attention to the limitations of current XAI methods.
  • We calculate each word’s positive and negative contribution score in a phishing email to show how it affects the phishing classification result.
  • We evaluate our method on a real-world dataset of phishing emails, showcasing the efficacy of the model and the clarity of the explanations it generates.
@@ 2 Literature Review
The field of cybersecurity has grown increasingly important and complex in the modern world. There are several subfields within cybersecurity, including data security, application security, network security, etc. Cyber-attacks of all kinds are becoming more and more frequent, and they can destroy or steal imported data. In addition, in this cybersecurity field, there is an enormous amount of new and creative techniques being developed daily for identifying and preventing cyberattacks. For the improvement of phishing email detection and prevention systems, many researchers have developed a wide range of novel strategies to counteract phishing emails, utilizing state-of-the-art methods from deep learning, machine learning, and hybrid models. Phishing is still a serious cybersecurity concern and one of the main ways hackers breach networks.
2.1 Classical Machine Learning Methods
@@
Detecting phishing attempts early on is critical for businesses nowadays. The security of an organization can be significantly increased by machine learning, which can identify potential assaults early on (Apruzzese et al., 2023). In ref. (Yasin and Abuhasan, 2016), a text processing, data mining, and knowledge discovery approach was combined to create an intelligent phishing email detection model. They utilized text stemming and WordNet ontology in pre-processing to enrich the model with synonyms and employed five machine learning classification algorithms (J48, Naïve Bayes, Support Vector Machine (SVM), Multi-Layer Perceptron (MLP), and Random Forest (RF)). They have notably achieved 99.1% accuracy with RF and 98.4% using J48 for an accredited dataset. The study (Harikrishnan et al., 2018) employed techniques such as SVD and NMF for feature extraction and dimensionality reduction from legitimate and phishing email text data. For the numeric representation of phishing emails, the author integrated TFIDF with the SVD and NMF, and Decision Trees, Logistic Regression, Random Forest, Naive Bayes, K-Nearest Neighbors (KNN), AdaBoost, and SVM have been utilized as classifiers. Hamid et al. (2013) presented a hybrid feature selection technique combining content-based and behaviour-based approaches to detect phishing emails. Identifying the behaviour-based features in phishing emails was the main objective of this research because an attacker cannot disguise these types of features. They examined the overall process with a phishing dataset that consists of 6923 emails, where phishing emails were collected from the Nazario dataset and legitimate emails from the SpamAssassin dataset, and achieved 94% accuracy with this hybrid feature selection technique. With an emphasis on the importance of top-ranking features, the study (Zamir et al., 2020) examines feature selection methods (GR, IG, Relief-F, RFE) on a dataset of 11,055 phishing sites with 32 attributes. To improve classifier accuracy, it suggests new attributes that are generated from the strongest and weakest features. A variety of supervised learning methods with differing feature sets are compared, such as RF, SVM, bagging, kNN, Neural Network (NN), and J48. The proposed method using RF showed 97.3%, and using stacking achieved 97.4% accuracy. Almomani et al. (2013) presented the Phishing Dynamic Evolving Neural Fuzzy Framework (PDENF), which detects and predicts unknown "zero-day" phishing emails with reduced false positive and false negative rates. It leverages hybrid and evolving connectionist systems for rapid, lifelong learning, featuring a minimal memory footprint and a simplified rule base for email classification. Salhi et al. (2021) demonstrated a novel email clustering approach that uses information gain techniques to evaluate email attributes, enabling effective differentiation between benign and harmful emails.
2.2 Deep Learning Methods
@@
In cybersecurity, deep learning has great potential, particularly for detecting network intrusions, classifying malware, and recognizing phishing attempts. Alhogail and Alsabih (2021) proposed a deep learning-based phishing email detection model based on graph convolutional network (GCN) and natural language processing (NLP). A publicly available balanced phishing dataset has been utilized, where the authors extracted features from the email body’s text to improve the proposed model’s accuracy. The model showed 98.2% accuracy with a low false-positive rate of 0.015%. For phishing detection, a new Intelligent Cuckoo Search (CS) Optimization Algorithm with a Deep Learning-based Phishing Email Detection and Classification (ICSOA-DLPEC) approach has been proposed in (Brindha et al., 2023). Data preprocessing (email cleaning, tokenization, and stop-word elimination) techniques were employed to derive the important features from the Enron email dataset. The Gated Recurrent Unit (GRU) model was used with a CS algorithm for the fine-tuning of the parameters of the proposed model and achieved a high accuracy of 99.72%. Dewis and Viana (2022) employed Long short-term memory (LSTM) model with NLP to detect both phishing and spam emails. They developed a phish responder, which is a Python-based command-line solution, and used deep learning and NLP for this classification task. The proposed LSTM model achieved 99% accuracy for the text-based dataset. Fang et al. (2019) introduced THEMIS, a phishing email detection model built on an improved recurrent convolutional neural network (RCNN) model, which was improved by using the Bidirectional Long Short-Term Memory (Bi-LSTM). 99.84% accuracy is demonstrated by the suggested THEMIS model on datasets that are not balanced, where the positive rate (FPR) is 0.043%.
In addition, researchers investigated deep neural architectures such as Recurrent Neural Networks (RNNs) and Deep Belief Networks (DBNs) (Zhang and Li, 2017; Bahnsen et al., 2017) in the context of phishing detection, as well as a Neural Network (NN) based on Reinforcement Learning (RL), called DENNuRL, for adaptive decision-making (Smadi et al., 2018).
2.3 Transformer-based Methods
@@
Moreover, transformer-based models are also making a good impact in the cybersecurity field.Thapa et al. (2023) concentrated on two specific models: RNN and BERT to test how well these models could identify phishing emails. They looked at how well these models performed when the email data was distributed unevenly across several organizations and when added more data. In this experiment, the BERT model’s test accuracy was 96.1%. In addition, another study (Atawneh and Aljehani, 2023) focused on different deep learning architectures (CNNs, LSTMs, RNNs, BERT) to identify the emails as legitimate or phishing by analyzing different features of emails. The proposed model using BERT and LSTM, reaching an impressive accuracy of 99.61% on publicly available 4 different datasets, outperforming other methods. Nowadays, other variants of the BERT model, such as DistilBERT, RoBERTa, and TinyBERT (Jiao et al., 2019), are performing well in the phishing and spam detection research sector (Jamal et al., 2023; Lee et al., 2020; Songailaitė et al., 2023). The authors (Jamal et al., 2023) presented an Improved Phishing and Spam Detection Model (IPSDM) which consists of two BERT variants, such as the fine-tuned DistilBERT and RoBERTA model,s and showed good accuracy in this research scope. Lee et al. Lee et al. (2020) proposed Context-Aware Tiny Bert (CatBERT), which exhibits faster processing and robustness to adversarial attacks than other transformer models, especially those that involve intentional keyword substitutions with synonyms or typos. The detection rate of this proposed model was 87%, which was far better than other compared models. Furthermore, researchers also used fine-tuned BERT-based models for phishing URL detection (Wang et al., 2023; Maneriker et al., 2021).
@@
Overall, this literature survey shows that although the existing studies demonstrate good performance using different datasets and models, most of them do not address model interpretability or explainability. Questions such as “How does the model work?” or “Which features contribute to phishing detection?” can arise in their methodologies. In this work, to answer those types of questions and solve the existing drawbacks in the literature, we have worked with explainable AI using not only one technique but also used two different techniques with different methodologies. Although several studies have conducted comprehensive experiments on phishing email detection, many of them relied on spam datasets rather than datasets specifically designed for phishing, which may limit the validity of their findings. In cybersecurity, we know that two different types of emails exist: spam and phishing, and not all spam emails are considered to be phishing. For this reason, we only focused on the phishing email dataset.
@@ 3 Methodology
In this paper, we have tried to explore the Transformer-based self-attention mechanism BERT model with explainable AI. Initially, we collected a dataset and prepared it with data cleaning and balancing techniques. After preparing the data, we have focused on optimizing and fine-tuning the model. For this phishing email detection purpose, we have selected the RoBERTa model based on our previous experimental analysis Uddin et al. (2025) and pre-trained it with the prepared data. The RoBERTa model demonstrates good performance in both balanced and imbalanced scenarios of datasets when identifying phishing emails from texts. At the end of the experiment, we employed XAI techniques to explore our fine-tuned model and described how it is working and performing with good accuracy in this text classification. The methodology of the proposed work is depicted in Fig. 1.
Figure 1: An overview of our experimental methodology.
3.1 Data Characteristics and Processing
For this experiment, a publicly available dataset has been collected from Kaggle. The dataset describes different kinds of email text bodies that can be used to identify phishing emails using in-depth text analysis. This dataset consists of two different types of email: Legitimate or Legitimate Email and Phishing Email. Fig. 2 shows the phishing email samples. The dataset comprises 11,322 legitimate Emails and 7,328 phishing Emails, indicating that it is imbalanced. Additionally, this dataset revealed the presence of null values in certain rows. As null values can’t be processed, that’s why those rows are removed.
We initially focused on data cleaning before addressing the data imbalance issue. This process involved converting the input text to lowercase, removing HTML tags that may be present in the text, replacing multiple whitespace characters with a single space, removing carriage return characters, and trimming leading and trailing whitespace. These steps ensured that the text was in a clear and consistent format for further processing.
After cleaning the data, we utilized a data augmentation technique called synonym replacement to address the imbalance problem. This technique replaces random words from the original text with their synonyms, preserving the original meaning of the sentence while maintaining its similarity to the original text. In this context, we randomly selected words that are not common stopwords and replaced them with their randomly chosen synonyms.
By using the augmentation technique, the minor class “Phishing Email” is oversampled. It is essential to address uneven data to develop equitable and precise predictive models. Balancing data ensures algorithms work well across all categories by reducing biases toward the majority class. This prevents skewed outputs and makes it possible to create a more accurate and equitable model for underrepresented classes. After oversampling, the dataset has 11,322 legitimate emails and 11,322 phishing emails in equal proportion. This data balancing also helps to reduce the risk of oversampling. Fig. 3 illustrates the imbalanced and balanced data before and after oversampling.
Figure 2: The sample data of Phishing Emails